[Brahms-l] Shift report 20040320 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sat Mar 20 2004 - 00:09:32 EST
------------------------------------------------------------
Date: Sat Mar 20 00:09:31 2004, Shift: 16:00-24:00
Supervisor: EBJ, with: BN
Log book#15 pages 91 - 91

Shift Summary:
------------------------------------------------------------
1. There was beam for most of the shift. 

2.  We took some data at the current settings after the DAQ issus were resolved.

3. Once the DAQ came back up, we didn't have any problems.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Mar 20 00:09:45 2004

This archive was generated by hypermail 2.1.8 : Sat Mar 20 2004 - 00:10:08 EST